This blog is to keep up-to-date info on the status of the UMass experiment.  The status of current projects should be updated every few days.   Information on broken or malfunctioning equipment should be kept here as well.  Any parts that need to be orderred should be listed as well.

12/3 Update

Where we're at:

Sample cell:  Thermocoax still needs to be installed.  Questions regarding this:  1)  Do we want to expose a long or short section of the thermocoax?  2)  How do we insulate the exposed section?, 3) Do we need to order some fine capillary?  We also need to work on the guts of the sample cell, i.e. circuit board, pin connections, etc.

Coax cable for fridge:  The first priority is to install the superconducting coax.  We need to decide on the length.  It needs to be long enough so that the SMA connector can be crimped on while in place in the fridge, but we don't want it to go too far up into the 1.5K space.  The cable needs to be cut to the precise specifications for crimping the connectors on it.  Yiming is going to contact Chao to find out these specifications.  The cable needs to be thermally cycled to LN2 several times to release any stresses.  The cable also needs a clamp at the 1.5K stage to prevent it from moving after installation.  Max has nearly perfected a technique of injecting Apiezon N grease into the displacer around the cable.  We should do a few more trial runs to be sure it works reliably.  We will then need to install the BeCu cable.  We decided to make the length such if the top reaches the top flange of the probe, the bottom end of the cable will be near the top of the 1.5K space.  It can then be pushed down to mate with the Nb coax, with attenuators, adapters and/or cold amplifiers in between.  The BeCu cable will need to be thermally anchored at this  stage using a Cu wire soldered near the attenuator.

Wiring:  We have Cinch connectors now.  We still need mounting hardware for the Cinch connectors.  The details of how to install the thermocoax into the fridge still have to be worked out.

Sample: Yiming is still working on a design for the sample.  The wire bonder in Yngvesson's lab is inaccessible while the lab is cleaned.  Yiming will talk to the students there to see when it might become available again.

Items to order

32 AWG teflon capillary(?)

 

Log in to read or post comments

Status update

Here's a current assessment of where we're at and what needs to be done.

Sample cell:  Max is done with machining and is currently putting the pieces together and testing.  He'll use copper tape to try and "plug the leaks" to figure out where the rf leaks are.  Thermocoax needs to be installed and the guts of the sample cell need to be put together.

Coax cable for fridge:  Copper coax still needs to be ordered.  We looked at the BeCu cable today.  It may ok for our needs.  Because of its high thermal conductivity, it will need to be thermally anchored at 1 K stage.  We will need to install attenuators at 1 K and mix chamber to filter out RT photons and noise.  We have five -30 dB SMA attenuators now. 

We still need to figure out how to install the superconducting cable in the displacer.  Yiming will contact Oxford and find out how we can install the coax:  Is packing it tightly in grease sufficient?  Does grease need to go up the entire hole or just down at the bottom?  What kind of grease?  Apiezon N?

Wiring:  Max is still looking into buying Cinch connectors, either from Oxford of directly from Cinch.  He will also try to figure out how to get the thermocoax installed given the tight space constraints.  For starters, he will see if he can fit a similar number of similar thickness wires into the fridge.

We still need mounting hardware for the Cinch connectors.

Sample: Yiming is still working on a design for the sample.

Log in to read or post comments

Projects underway

Sample cell:  Max is working on machining this. 

Coax cable for fridge:  Yiming is going to order SS cable to run from top of probe to 1K pot.  Some copper coax will also be ordered.  Connectors will need to be attached to the cables as they are installed.  We will need to figure out how to install a cable in the displacer without introducing any leak.  We may need to buy or machine an adaptor that surrounds the cable and fits snuggly into the displacer.

Wiring:  We will use existing wiring on the fridge.  Max is looking into buying Cinch connectors, either from Oxford of directly from Cinch.

Filter box:  Max was going to measure the reponse of these filters, which appear to be pi-configured Butterworth filters with L = 1 mH and C = 3 nF.  A quick calculation suggests that the filter should have a 3dB point of ~130 kHz and an impedance of 408 ohms.  This needs to be confirmed.

Thermocoax:  We need to figure out how to install this in the fridge within the tight space constraints.  The cables will likely need to pass up and down the mixing chamber several times through the grooves in G10.  The center conductors of the cables will need to be soldered onto the Cinch connector.  Some simple mounting hardware should be made so that the Cinch connectors are not simply dangling.

Sample: Yiming is working on a design for the sample.  Once the evaporator is working, we should try to make several versions of the device for testing.

Log in to read or post comments

Malfunctioning equipment

At this point, the evaporator at UMass won't pump down below 10^-4 Torr.  According to Stefan, the problem persists when the gate valve is closed.  No problem was found with the foreline valve.  It is likely that there is something wrong with the diffusion pump at this point.  It may have to be cleaned.